# Are LLMs Naturally Good at Synthetic Tabular Data Generation?

### v2

@@
Abstract
Large language models (LLMs) have demonstrated their prowess in generating synthetic text and images; however, their potential for generating tabular data—arguably the most common data type in business and scientific applications—is largely underexplored. This paper demonstrates that LLMs, used as-is, or after traditional fine-tuning, are severely inadequate as synthetic table generators. Due to the autoregressive nature of LLMs, fine-tuning with random order permutation runs counter to the importance of modeling functional dependencies, and renders LLMs unable to model conditional mixtures of distributions (key to capturing real world constraints). We showcase how LLMs can be made to overcome some of these deficiencies by making them permutation-aware. Our code and data are anonymously hosted [^1] https://github.com/ShengzheXu/Permutation-aided-Fine-tuning.
1 Introduction
@@
Large language models (LLMs) have found applicability in a rich diversity of domains, far beyond their original roots [1, 4, 22, 25, 38, 31, 33]. As so-called foundation models [18] they have been shown to be re-targetable to a variety of downstream tasks. Our focus here is to view LLMs as raw synthetic tabular data generators rather than as supporting an analysis or discovery task. Arguably, LLMs are adept at synthetic generation of text, images, videos, code, documentation, and many other modalities. Unlike these traditional modalities, the generation of tabular data requires the generative mechanism to effectively estimate mixtures of distributions, to correctly capture relationships between columns in tables.
Related Work: In fact, the synthetic generation of this type of data has been extensively studied and continues to be a strong focus of machine learning research. Various approaches have been developed for tabular data generation: Lei et al. [40] proposed CTGAN where rows are independent of each other; a conditional GAN architecture ensures that the dependency between columns is learned. Tabsyn [43] also generate independent rows but with a VAE plus diffusion approach. DoppelGanger [19] uses a combination of an RNN and a GAN to incorporate temporal dependencies across rows but this method has been tested in traditional, low-volume settings such as Wikipedia daily visit counts. For high-volume applications, STAN [42] utilizes a combination of a CNN and Gaussian mixture neural networks to generate synthetic network traffic data. GraphDF  [6] conducts multi-dimensional time series forecasting. GOGGLE  [20] employs a generative modeling method for tabular data by learning relational structures.
However, the use of language models (LLMs) for tabular data generation is still underexplored. Most modern LLMs are based on the transformer architecture  [39] with parameters ranging from few millions to billions  [15], and researchers have developed creative ways to harness LLMs in traditional machine learning and data contexts. LIFT [11] initially transforms a table row into a sentence, such as ‘An Iris plant with sepal length 5.1cm, sepal width 3.5cm’, and employs an LLM as a learning model for table classification, regression, and generation tasks. GReaT [3] utilizes a GPT-2 model that has been fine-tuned using a specific corpus for synthetic data generation. They also show that even small-scale models such as Distill-GPT [30] have the potential for synthetic data generation [3]. These models are specially viable for tabular generation given the lower compute costs of aligning smaller models to large and varied tabular datasets. A general benefit of utilizing LLMs is the promise of eliminating customized preprocessing pipelines. Following this paradigm, in this paper, we inspect just how well LLMs fare at generating traditional tabular data, the lingua franca of machine learning.
@@
The ‘impedance mismatch’ between autogressive LLMs and synthetic data generation. The most popular incarnations of language models are auto-regressive models, e.g., LLama [38], GPT-x [25]. Thus each word or token is generated conditional on past tokens in a sequential manner using attention models. In a synthetic data context, each ‘sentence’ typically represents a row of tabular data, and each ‘word’ corresponds to an attribute in that row. The previous state-of-the-art models (GReaT [3]), has advocated the use of random feature orders but as we show in Fig. 1, when fine-tuning is done with random feature orders, key relationships are often not captured or, worse, violated. In particular, with tabular data, there are numerous functional dependencies at play (e.g., the applicable range of latitudes and longitudes are constrained given a state) and as a result, generating tokens in random orders is bound to cause violations.
Regrettably, the underlying problem is not just with the methods used for synthetic data generation but also the measures used to evaluate results. Popular contemporary evaluation criteria might overlook whether the synthetic data faithfully reproduces (or violates) underlying inherent functional dependencies. For instance, synthetically generated data might be able to pass a test on an underlying variable’s distribution or even a test such as a Pearson coefficient with only two variables. For multiple variables, measures such as machine learning efficiency (MLE) [3, 40, 43, 11] are frequently used to construct classifiers or regressors but it is possible for the resulting models to operate effectively without capturing key interdependence or functional relationships. Further complexities abound due to the multitude of data types in tables, e.g., relationships between numerical and categorical variables.
In light of these challenges, we present our work to investigate in depth, how the unique characteristics of tabular data manifest as challenges in an LLM-based generative context and propose an effective solution. The main contributions of this paper are:
@@
$\bullet$ We highlight an important deficiency with using LLMs for synthetic tabular data generation and explore the performance of many state-of-the-art generation models in the context of composite and multi-category tabular schema.
@@
$\bullet$ We inject knowledge of pre-existing functional relationships among columns into the autoregressive generation process, so that the generated synthetic data respect more real constraints. In particular, We present a taxonomy of functional dependencies (FDs) whose discovery and organization into a column dependence graph supports their incorporation into the LLM fine-tuning process via a permutation function, leading to our approach dubbed Permutation-aided Fine-tuning.(PAFT ).
@@
$\bullet$ We evaluate the performance of PAFT on a range of datasets featuring a diverse mix of attribute types, functional dependencies, and complex relationships. Our results demonstrate that PAFT is the state-of-the-art in reproducing underlying relationships in generated synthetic data.
@@
$\bullet$ Finally, we demonstrate through rigorous experiments that relying just on standard univariate distribution, bivariate correlation, and even the evaluation of downstream machine learning models (which primarily focuses on predicting a single column in a dataset) is grossly insufficient for assessing the quality of synthetic data and propose systematic remedies like measuring violation rate of known domain rules.
2 Challenges to Synthetic Table Generation in the Current LLM Paradigm
Figure 1: Synthetic location data generated by GReaT for the state of Delaware (DE) in the United States, where the solid line represents the official border of Delaware and the blue and yellow colors indicate the legal and illegal samples in the synthetic data respectively. Traditional GReaT fine-tuning with random order permutation (Fig 1 b) works well for one state at a time but creates datasets with egregious violations when applied to all 51 states (Fig 1 c). PAFT, our fine-tuning approach that utilizes an established optimal fine-tuning order, significantly mitigates the challenges posed by this complex dataset (Fig 1 d). 51 states includes (50 US states + Washington DC).
In the field of tabular data generation, generative models that focus on conditional distribution learning (auto-regressive models [42] and LLM [3]) have a distinct advantage over generative models based on joint distribution learning (GAN [40], VAE [40, 43], and Diffusion [43]). These models excel at accurately generating conditional constraints between columns, both in theory and in practice [11, 3]. However, the performance of existing conditional generation techniques can differ depending on the dataset, and even within different categories of the same dataset. It is worth noting that this paper analyzes conditional constraints between various types of attributes, rather than just focusing on numerical correlations. These conditional constraints often exhibit asymmetry, which has the potential to greatly influence the training and generation process of the model.
2.1 Can LLMs model composite datasets?
By a composite dataset, we mean one that is best modeled as a mixture of distributions rather than a single distribution for the full table. The traditional approach to fine-tuning pioneered by methods such as GReaT work only if we assume that the table can be modeled by a single distribution rather than a mixture. For instance, in a geolocation dataset, fine-tuning with GReaT works for one US state at a time.
Figure 2: For a composite dataset, comparison of state-specific violation rates for different synthetic data generation approaches. The states (x-axis) are sorted based on increasing violations. PAFT significantly mitigates state-specific violations in a composite dataset.
In Fig. 1, we aim to generate synthetic location data made up of (latitude, longitude) tuples for US states. Fig. 1 (a) depicts sampled real locations for the state of Delaware (DE) chosen for its irregular shape. In Fig. 1 (b), we depict the results from GReaT which shows a very admirable 1% violation rate. Following their approach, we fine-tune the Distill-GPT2 LLM [35] by supplying tuples of (latitude, longitude, state) in a randomized order.
However, when we extend this method to all 51 US states and territories, we can see that the generated data for Delaware drops significantly in quality, with violations growing to 23% after fine-tuning (Fig. 1 (c)). Generated data for other states suffer from similar violations. With our fine-tuning approach (PAFT) presented later (Fig. 1 (d) we are able to support the generation of synthetic data for all 51 states with minimal violations.
Observation 1: Modeling one class can be done in an order-agnostic manner but modeling a multi-class dataset requires paying careful attention to feature ordering.
Observation 2:. Within a composite, multi-class, dataset, there is variance in difficulty of modeling specific classes (e.g., New Mexico is easier to model for LLMs than Delaware or Alaska). See Fig. 5 and Fig. 10 in Appendix.
2.2 How does the importance of feature order vary with schema complexity?
Observation 3: While any order can achieve satisfactory results for unmixed data, it becomes increasingly crucial to employ a better order when the dataset becomes progressively more composite. We conduct a simple experiment based on a simulated table with four attributes and corresponding six column-wise dependencies (see Appendix C.1). Figure 3 shows the relationship between different order permutations. In this simulated table, there are a total of six dependencies between columns (to be made precise using the notion of FDs later); see Appendix C.1. The evaluation measure captures the extent to which the generated data preserves the functional dependencies present in the original data, so that 6 means the generated data is faithful to the original schema complexity and 0 means the data has maximum violations. From Figure 3, we observe that as the schema complexity increases (progressively adding more classes from left to right), it becomes increasingly important to establish a precise sequence in order to provide higher quality generated data. Thus, with randomly guessed order permutations, auto-regressive generation with LLMs are severely limited in their ability when it comes to simulating composite datasets with complex functional dependencies.
2.3 How susceptible are LLMs and other models to sacrificing conditional nuance in favor of capturing joint distributions?
Figure 3: For a simulated table with a total of 6 functional dependencies (FDs), we showcase increased FD violations as the dataset becomes progressively more composite (increasing classes). x-axis represents degree of deviation of data permutation from ideal permutation while y-axis represents number of retained FDs (higher is better, with FD = 6 being optimal). Each blue dot in a graph represents a single possible permutation of the data in fine-tuning process. Thus, while order permutation may not be as crucial for unmixed datasets (left), permutation order becomes important as dataset complexity increases (right).
A recently presented paper, TabSyn [43], showcased remarkable advancements in joint-distribution learning generative models, surpassing previous models of similar lineage, in terms of distributional correlation measures and machine learning efficiency. However, as we will show, methods like TabSyn fail to capture real-world dependencies, e.g., the relationship between protocol and port number in network traffic data [42], or the geographical boundaries of US states as described earlier. Constraints in tabular data can be inferred empirically from conditional distributions in the data and thus methods such as LLMs are particularly powerful because of their autoregressive nature.
Observation 4: While joint distribution learning models (CTGAN [40], CorpulaGAN [40], TabSyn [43]) and conditional distribution models (GReaT and ours) exhibit comparable performance in modeling joint distributions, LLM-based fine-tuning demonstrates greater advantages for capturing conditional distributions, functional relationships, and domain constraints (see Figure 5 and Table 2).
@@ 2.2 How does the importance of feature order vary with schema complexity?
3 PAFT : Permutation-Aided Fine-Tuning
@@
3.1 Tabular Data Generation with LLMs
While training $p_{\theta}(\cdot)$, it is usually assumed that all records $\mathbf{x}_{i}\in\mathcal{D}$ are independent. Generating new data samples $\hat{\mathbf{x}}_{i}\in\hat{\mathcal{D}}$ can be done in various ways (e.g., see [41, 40, 45]) which aim to directly estimate the joint distribution $\mathbb{P}(\mathcal{A})$ or, as is done here in PAFT , where $\mathbb{P}(\mathcal{A})$ is estimated by an autoregressive LLM based generative process, as a product of multiple conditional densities governed by the input ordering.
@@ 3.1 Tabular Data Generation with LLMs
$\mathcal{L}(\theta;\mathcal{D})=-\sum\limits_{\mathbf{x}_{i}\in\mathcal{D}}\sum\limits_{j=1}^{w}\mathrm{log}\,\mathbb{P}(x_{i}|\mathbf{x}_{(i,1:j-1)})$ (1)
@@
$\mathbb{P}(\textbf{x}_{i})=\mathbb{P}(x_{i,1},...,x_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(x_{i,j}|x_{i,1},..,x_{i,j-1})$ (2)
@@
$\begin{split}t_{i,j}=[c_{j},\textrm{`}is\textrm{'},x_{i,j},\textrm{`},\textrm{'}],&i\in\{1,..,n\},j\in\{1,..,m\},\\ \mathbf{t}_{i}=[t_{i,1},t_{i,2},..,t_{i,m}],&i\in\{1,..,n\}\end{split}$ (3)
@@
Since we consider autoregressive LLM-based generative models, employing the chain rule to sequentially produce each column of a table record $\mathbf{t}_{i}$, we can view each generation step as approximating the joint distribution of the table columns as a product of conditional distributions (i.e., $\mathbb{P}(t_{i,1},...,t_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,j}|t_{i,1},..,t_{i,{j-1}})$). However, as the number of columns increases and the relationships between columns get more conditional, the likelihood of encountering training and sampling bias due to class imbalance also rises [40]. To minimize such adverse effects, we can consider injecting knowledge of the pre-existing functional relationships among columns, to govern the autoregressive generation process. To infer such functional relationships, we leverage a learned dependency graph derived from functional dependency (FD) relations which enables us to effectively determine the appropriate training and sampling sequence. This, in turn, allows us to alleviate potential biases during training by establishing a generation curriculum leading to improved estimation accuracy of the joint distribution $\mathbb{P}(\textbf{t}_{i})$ in auto-regressive prediction $\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{m}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}})$, where the ordering $t_{i,k_{1}}\dots,t_{i,k_{m}}$ is obtained by a feature order permutation function $\mathbf{k}=\phi(\mathcal{D},\mathcal{S})$. We detail the requisite background and design of $\phi(\mathcal{D},\mathcal{S})$ in sections 3.2 and  3.3.
3.2 Discovery and Distillation of Functional Dependencies (FD)
A functional dependency (FD) is a relationship $R$ in schema $S$ that exists when a subset of attributes $A\subset S$ uniquely determines another subset $B\subset S$ of attributes. We succinctly represent an FD as $R:A\rightarrow B$ which specifies that $B$ is functionally dependent on $A$. FDs can be record-level or schema-level.
Definition 1 (Record-Level FD).
With $A$, $B$ being two disjoint subsets of the schema (columns) of table $\mathcal{D}$, let $a\in A$, $b\in B$ denote two entries in A and B; a record-level FD f has the form: $f:a\rightarrow b$.

### v3

# Why LLMs Are Bad at Synthetic Table Generation

@@
Abstract.
Synthetic data generation is integral to ML pipelines, e.g., to augment training data, replace sensitive information, and even to power advanced platforms like DeepSeek. While LLMs fine-tuned for synthetic data generation are gaining traction, synthetic table generation—a critical data type in business and science—remains under-explored compared to text and image synthesis. This paper shows that LLMs, whether used as-is or after traditional fine-tuning, are inadequate for generating synthetic tables. Their autoregressive nature, combined with random order permutation during fine-tuning, hampers the modeling of functional dependencies and prevents capturing conditional mixtures of distributions essential for real-world constraints. We demonstrate that making LLMs permutation-aware can mitigate these issues. Our code and data are anonymously hosted [^1] https://github.com/ShengzheXu/Permutation-aided-Fine-tuning.
copyright: acmlicensedjournalyear: 2025
1. Introduction
@@
Large language models (LLMs) have found applicability in a rich diversity of domains, well beyond their original roots (Achiam et al., 2023; Brown et al., 2020; Miller and Blair, 2009; OpenAI, 2022; Touvron et al., 2023; Raffel et al., 2020; Romera-Paredes et al., 2023). As so-called foundation models (Liang et al., 2022) they have been shown to be re-targetable to a variety of downstream tasks. Our focus here is to view LLMs as raw synthetic tabular data generators rather than as supporting an analysis or discovery task. Arguably, LLMs are adept at synthetic generation of text, images, videos, code, documentation, and many other modalities. The use of LLMs to generate tabular data is quite understudied.
@@
Such synthetic tabular data generation is integral to ML pipelines, e.g., too augment training data, replace sensitive information, and even to power advanced platforms like DeepSeek (Liu et al., 2024; Guo et al., 2025). However, the unique characteristics of tabular data manifest as challenges in an LLM-based generative context. The most popular incarnations of language models are auto-regressive models, e.g., LLama (Touvron et al., 2023), GPT-x (OpenAI, 2022), DeepSeek (Liu et al., 2024; Guo et al., 2025), wherein each word or token is generated conditional on past tokens in a sequential manner using attention models. In a synthetic data context, each ‘sentence’ typically represents a row of tabular data, and each ‘word’ corresponds to an attribute in that row. The previous state-of-the-art models (GReaT (Borisov et al., 2022)), has advocated the use of random feature orders but as we show in Table 1, when fine-tuning is done with random feature orders, key relationships are often not captured or, worse, violated. In particular, with tabular data, there are numerous functional dependencies at play and as a result, generating tokens in random orders is bound to cause violations. There is thus an ‘impedance mismatch’ between autogressive LLMs and synthetic data generation.
The main contributions of this paper are:
@@
  • We highlight an important deficiency with using LLMs for synthetic tabular data generation and explore the performance of many state-of-the-art generation models in the context of composite and multi-category tabular schema.
@@
  • We inject knowledge of pre-existing functional relationships among columns into the autoregressive generation process, so that the generated synthetic data respects more real constraints. In particular, we present a taxonomy of functional dependencies (FDs) whose discovery and organization into a column dependence graph supports their incorporation into the LLM fine-tuning process via a permutation function, leading to our approach dubbed Permutation-aided Fine-tuning.(PAFT ).
@@
  • We evaluate the performance of PAFT on a range of datasets featuring a diverse mix of attribute types, functional dependencies, and complex relationships. Our results demonstrate that PAFT is the state-of-the-art in reproducing underlying relationships in generated synthetic data.
@@
  • Finally, we demonstrate through rigorous experiments that relying just on standard univariate distribution, bivariate correlation, and even the evaluation of downstream machine learning models (which primarily focuses on predicting a single column in a dataset) is grossly insufficient for assessing the quality of synthetic data and propose systematic remedies like measuring violation rates of known domain rules.
2. Related Work
Table 1. Comparison of multiple synthetic data generation approaches. The second row showcases the state of West Virginia (WV), which is a subset of the whole data. In the figures, the solid line represents the official border of West Virginia and the Blue and Orange colors indicate the legal and illegal samples in the synthetic data respectively.
Whole Data
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Sub-Category
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Real CTGAN CopulaGAN TabSyn (Diffusion) GReaT (LLM-GPT) PAFT (LLM-GPT)
Tabular data synthesis and representation learning for tables have been extensively studied (Du et al., 2021; Wang et al., 2021; Zhang et al., 2023b; Margeloiu et al., [n. d.]; Fang et al., 2024; Patki et al., 2016; Du and Li, 2024; McKenna et al., 2019; Park et al., 2018; Wang et al., 2020). One key theme has been the discovery of functional dependencies (FDs) which has been studied by the data mining community from both theoretical and application points of view (Zhang et al., 2020; Chen et al., 2019; Muralidhar et al., 2018; Mandros et al., 2020; Pennerath et al., 2020; Zheng et al., 2023). We carve out related work into two sections, for our purposes: pre-LLM (or non-LLM) approaches for synthetic table generation which continue to hold the mainstay, and LLM approaches.
Lei et al. (Xu et al., 2019) proposed CTGAN where rows are independent of each other; a conditional GAN architecture ensures that the dependency between columns is learned. Tabsyn (Zhang et al., 2023b) showcased remarkable advancements in joint-distribution learning via a VAE plus diffusion approach, surpassing previous models of similar lineage, in terms of distributional correlation measures and machine learning efficiency. DoppelGanger (Lin et al., 2020) uses a combination of an RNN and a GAN to incorporate temporal dependencies across rows but this method has been tested in traditional, low-volume settings such as Wikipedia daily visit counts. For high-volume applications, STAN (Xu et al., 2021) utilizes a combination of a CNN and Gaussian mixture neural networks to generate synthetic network traffic data. GraphDF  (Chen et al., 2023) conducts multi-dimensional time series forecasting. GOGGLE  (Liu et al., 2022) employs a generative modeling method for tabular data by learning relational structures.
The use of language models (LLMs) for tabular data generation is still underexplored. Most modern LLMs are based on the transformer architecture  (Vaswani et al., 2017) with parameters ranging from few millions to billions  (Hoffmann et al., 2022), and researchers have developed creative ways to harness LLMs in traditional machine learning and data contexts. LIFT (Dinh et al., 2022) initially transforms a table row into a sentence, such as ‘An Iris plant with sepal length 5.1cm, sepal width 3.5cm’, and employs an LLM as a learning model for table classification, regression, and generation tasks. GReaT (Borisov et al., 2022) utilizes a GPT-2 model that has been fine-tuned using a specific corpus for synthetic data generation. They also show that even small-scale models such as Distill-GPT (Radford et al., 2019) have the potential for synthetic data generation (Borisov et al., 2022). These models are specially viable for tabular generation given the lower compute costs of aligning smaller models to large and varied tabular datasets. A general benefit of utilizing LLMs is the promise of eliminating customized preprocessing pipelines.
A theme that will be pertinent to our work is the idea of feature ordering (FO) for tabular data generation which has been investigated from multiple angles (Sharma et al., 2024; Distiawan et al., 2018; Ribeiro et al., 2019; Zhu et al., 2022). There are also several approaches (e.g., (Kanter and Veeramachaneni, 2015; Cvitkovic, 2020)) that synthesize, discover, or aggregate features from relational databases, leveraging order information when possible, for use in machine learning pipelines. It worth noting that even in the LLM community, the task of context sorting for LLM prompting is not trivial and has gained significant attention lately (Chen et al., 2024).
3. Challenges to Synthetic Table Generation in the Current LLM Paradigm
Figure 1. Overview of the proposed Permutation-Aided Fine-tuning (PAFT ) approach.
Admittedly, the sure-fire way to check if a dataset has been faithfully modeled is to see if the joint distribution of its features is captured accurately. Most current tests of synthetic data generation quality focus on fidelity to single-column distributions, or to multi-column distributions. For multiple variables, measures such as machine learning efficiency (MLE) (Borisov et al., 2022; Xu et al., 2019; Zhang et al., 2023b; Dinh et al., 2022) are frequently used to construct classifiers or regressors.
A key lesson from probabilistic graphical model research (Koller and Friedman, 2009) is that factorizing joint distributions into products of conditionals (e.g., Bayesian networks) dramatically helps reduce the number of parameters necessary to capture the underlying data characteristics. A similar lesson from database research (Garcia-Molina et al., 2008) is that modeling functional dependencies (FDs) in data helps reduce redundancy in modeling and storage. These lessons, i.e., that order matters, continue to apply in the LLM era but are not internalized in our prompt ordering, fine-tuning, or evaluation methodologies. Other researchers have noted the importance of ordering in LLMs (Chen et al., 2024; Prystawski et al., 2024) but this lesson has not been leveraged to improve synthetic table generation by LLMs. In the absence of leveraging good feature orders, existing approaches either focus on ‘one order’, ‘no order’, or ‘all orders’. All of these approaches severely limit the quality of generated synthetic data.
As an example, Table 1 presents a case study on using models to generate synthetic data of locations in various states of the USA. The data is of the form (state, latitude, longitude) where the attributes adhere to the FD: $\{\text{latitude},\text{longitude}\}\rightarrow\text{state}$. Existing methods can generate satisfactory univariate distributions (as shown in the border of the top row plots) but fail to capture the joint distribution (center of the top row plots) and conditional distributions across subcategories (bottom row plots).
@@ 3. Challenges to Synthetic Table Generation in the Current LLM Paradigm
Figure 1 shows an overview of the proposed permutation aided fine-tuning approach (PAFT ). A typical workflow is 1) Textual Encoding (Section 4.1) 2) Functional Dependency (FD) discovery (Section 4.2) 3) FD Distillation ( Section 4.2) and 4) Feature Order Permutation Optimization ( Section 4.3). Our fine-tuning and sampling strategy is explained in Section 4.4.
4. PAFT : Permutation-Aided Fine-Tuning
@@
4.1. Tabular Data Generation with LLMs
While training $p_{\theta}(\cdot)$, it is usually assumed that all records $\mathbf{x}_{i}\in\mathcal{D}$ are independent. Generating new data samples $\hat{\mathbf{x}}_{i}\in\hat{\mathcal{D}}$ can be done in various ways (e.g., see (Xu and Veeramachaneni, 2018; Xu et al., 2019; Zhao et al., 2021)) which aim to directly estimate the joint distribution $\mathbb{P}(\mathcal{A})$ or, as is done here in PAFT , where $\mathbb{P}(\mathcal{A})$ is estimated by an autoregressive LLM based generative process, as a product of multiple conditional densities governed by the input ordering.
@@ 4.1. Tabular Data Generation with LLMs
(1) $\mathcal{L}(\theta;\mathcal{D})=-\sum\limits_{\mathbf{x}_{i}\in\mathcal{D}}\sum\limits_{j=1}^{w}\mathrm{log}\,\mathbb{P}(x_{i}|\mathbf{x}_{(i,1:j-1)}).$
@@
(2) $\mathbb{P}(\textbf{x}_{i})=\mathbb{P}(x_{i,1},...,x_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(x_{i,j}|x_{i,1},..,x_{i,j-1}).$
@@
(3) $\begin{split}t_{i,j}=[c_{j},\textrm{`}is\textrm{'},x_{i,j},\textrm{`},\textrm{'}],&i\in\{1,..,n\},j\in\{1,..,m\},\\ \mathbf{t}_{i}=[t_{i,1},t_{i,2},..,t_{i,m}],&i\in\{1,..,n\}.\end{split}$
@@
Since we consider autoregressive LLM-based generative models, employing the chain rule to sequentially produce each column of a table record $\mathbf{t}_{i}$, we can view each generation step as approximating the joint distribution of the table columns as a product of conditional distributions (i.e., $\mathbb{P}(t_{i,1},...,t_{i,m})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,j}|t_{i,1},..,t_{i,{j-1}})$). However, as the number of columns increases and the relationships between columns get more conditional, the likelihood of encountering training and sampling bias due to class imbalance also rises (Xu et al., 2019). To minimize such adverse effects, we can consider injecting knowledge of the pre-existing functional relationships among columns, to govern the autoregressive generation process. To infer such functional relationships, we leverage a learned dependency graph derived from functional dependency (FD) relations which enables us to effectively determine the appropriate training and sampling sequence. This, in turn, allows us to alleviate potential biases during training by establishing a generation curriculum leading to improved estimation accuracy of the joint distribution $\mathbb{P}(\textbf{t}_{i})$ in auto-regressive prediction $\mathbb{P}(t_{i,k_{1}},...,t_{i,{k_{m}}})\simeq\prod_{j=1}^{m}\mathbb{P}(t_{i,k_{j}}|t_{i,k_{1}},..,t_{i,k_{j-1}})$, where the ordering $t_{i,k_{1}}\dots,t_{i,k_{m}}$ is obtained by a feature order permutation function $\mathbf{k}=\phi(\mathcal{D},\mathcal{S})$. We detail the requisite background and design of $\phi(\mathcal{D},\mathcal{S})$ in sections 4.2 and  4.3.
4.2. Discovery and Distillation of Functional Dependencies (FD)
A functional dependency (FD) is a relationship $R$ in schema $S$ that exists when a subset of attributes $A\subset S$ uniquely determines another subset $B\subset S$ of attributes. We succinctly represent an FD as $R:A\rightarrow B$ which specifies that $B$ is functionally dependent on $A$.